# Thinking in Cocktail Party: Chain-of-Thought and Reinforcement Learning for Target Speaker Automatic Speech Recognition

### v1

@@ 3.2 Experiment settings
The proposed model uses Qwen2.5-0.5B-Instruct [25] as the backbone LLM. It adopts a pre-trained Data2Vec2 model as the audio encoder, which is trained on one million hours of unsupervised speech data. The number of parameters of Data2Vec2 is 0.3 billion. During the SFT stage, the AdamW optimizer [26] is applied with a learning rate schedule combining linear warmup and cosine annealing. The learning rate is gradually increased from 1e-7 to 7e-5 over the first 3000 steps and decayed to a minimum of 1e-6. The model is trained for 240000 steps, equivalent to one full epoch, with a global batch size of 112. To prevent input truncation, the maximum text length is set to 512. For RL training, the model is fine-tuned on 20000 selected samples for 2000 steps with a learning rate of 1e-6 and a temperature of 1. For each input, 8 responses are generated, with a maximum length of 512 tokens.
@@ 3.3 Evaluation metrics
In this experiment, the evaluation metric is WER, which is calculated by extracting the text between <answer>and </answer>tags from the decoded output of the LibriSpeech, Libri2Mix, and Libri3Mix test sets. For the small number of outputs with formatting errors or incomplete tags, the result is treated as empty for evaluation. To reflect overall performance, a weighted average is computed with the sample sizes as weights.
@@ 4.1 Comparison to baselines
We compare our proposed model with both traditional TS-ASR methods and LLM-based approaches. As shown in Table 1, our model achieves significant improvements, with WERs of 4.79% on Libri2Mix and 12.23% on Libri3Mix.
@@
Compared with MT-LLM, the current best-performing LLM-based TS-ASR model, our TS-ASR BASE model without CoT data has a lower performance. However, training with CoT data achieves a significant improvement, with the average WER relatively reduced by 26.18%. It demonstrates that logic reasoning helps the model more accurately transcribe the target speaker. Based on the model trained on CoT data, RL fine-tuning is applied on selected data, further reducing the average WER to 8.33%. It indicates that RL training can help the model further improve the reasoning capabilities. Our model outperforms all baseline models in average WER, achieving new SOTA performance.
Table 1: TS-ASR performance on LibriSpeech, Libri2Mix and Libri3Mix test-clean dataset, as well as their weighted average. Evaluated by WER(%).
@@
Single Libri2Mix Libri3Mix Avg
WavLM + TSE [9] - 7.6 - -
TS-VAD(460h)  [12] - 6.61 14.81 -
MT-LLM  [13] 2.3 6.7 16.2 11.09
TS-ASR Base 8.1 7.4 17.24 12.45
TS-ASR + CoT 4.08 5.21 13.48 9.19
TS-ASR + CoT + RL 3.47 4.81 12.23 8.33
4.2 Ablation studies for CoT
We explore the impact of different CoT data and training strategies on model performance to identify an effective configuration. As shown in Table 2, the model trained on the data using continuous similarity scores is compared with that with discrete similarity levels. The results show that the use of discrete similarity levels outperforms using continuous similarity scores, which indicates that replacing continuous floating values with discrete levels in CoT data leads to more stable and effective model convergence. The benefit is particularly evident in multi-speaker scenarios, indicating that similarity information indeed helps the model identify the target speaker from overlapping speech, which is consistent with our expectations. In addition, the proposed random reasoning strategy mentioned in 2.2.2 is implemented. Experimental results show that this approach obtains a marginal improvement compared with using full CoT data, especially on Libri2Mix and Libri3Mix.
Table 2: Ablation Studies for CoT, evaluated by WER(%). FC indicate the use of Full CoT data and $R$ denotes the random reasoning.
Model LibriSpeech
Single Libri2Mix Libri3Mix Avg
TS-ASR + CoT
FC + similarity score 3.8 5.39 14.09 9.52
FC + similarity level 3.88 5.29 13.57 9.23
RR + similarity level 4.08 5.21 13.48 9.19
4.3 Ablation studies for RL
Table 3 presents the impact of different training data selection strategies for RL on model performance. Specifically, “Random Sampling” denotes randomly selecting a subset from the full training dataset for RL training. “Balanced Correct&Error” selects correct and error samples in a 1:5 ratio from the model trained on CoT data for RL training, aiming to focus on difficult cases while preserving general performance. The ”Error-only Sampling” strategy uses only incorrectly predicted samples for subsequent training.
Experimental results show that the “Random Sampling” strategy performs worst, with an average WER of 8.73%, while the “Error-only Sampling” strategy achieves the best result, reducing WER to 8.33%. This demonstrates that in reinforcement learning for TS-ASR, challenging error samples contribute more to performance improvement than correct samples. Therefore, training with a higher proportion of error samples is more effective for helping the model correct mistakes and improve overall accuracy.
Table 3: Ablation Studies for RL, evaluated by WER(%). $R$ indicate the random reasoning strategy.
Model LibriSpeech
Single Libri2Mix Libri3Mix Avg
TS-ASR + CoT (with RR + similarity level) + RL
+ Random Sampling 4.08 5.01 12.70 8.73
+ Balanced Correct&Error 3.91 4.8 12.39 8.47
+ Error-only Sampling 3.47 4.81 12.23 8.33
@@ 5 Conclusion
In this work, we proposed a novel framework that incorporates CoT and RL training into the TS-ASR task to enhance reasoning capabilities in cocktail party scenarios. The approach began with training a TS-ASR BASE model using the LALM architecture. Subsequently, a novel methodology was proposed to generate CoT training samples by extracting salient features from overlapped speech, such as number of speakers, speech duration, gender information, and similarity to the reference speech. These features were systematically organized into structured reasoning data to fine-tune the TS-ASR BASE model. Furthermore, a subset of mispredicted samples from the CoT training phase was randomly selected for additional RL-based refinement. Experimental results showed a significant improvement of TS-ASR performance with CoT and RL training, which not only demonstrated the effectiveness of our constructed CoT dataset and the proposed training framework but also established a SOTA performance for TS-ASR on comparable benchmark datasets. The successful integration of CoT and RL into TS-ASR demonstrated the efficacy of logical reasoning in cocktail party scenarios. As LALMs continue to advance with CoT and RL techniques, we believe this framework has significant potential for further solving the cocktail party problem.

### v2

@@ 3.2 Experiment Settings
The proposed model uses Qwen2.5-0.5B-Instruct (Team, 2024) as the backbone LLM. It adopts a pre-trained Data2Vec2 model as the audio encoder, which is trained on one million hours of unsupervised speech data. The number of parameters of Data2Vec2 is 0.3 billion.
During the SFT stage, the AdamW optimizer (Loshchilov and Hutter, 2019) is applied with a learning rate schedule combining linear warmup and cosine annealing. The learning rate is gradually increased from 1e-7 to 7e-5 over the first 3000 steps and decayed to a minimum of 1e-6. The model is trained for 240000 steps, equivalent to one full epoch, with a global batch size of 112. To prevent input truncation, the maximum text length is set to 512. Finally, the model with the best performance on the validation set is selected as the final model.
For RL training, the model is fine-tuned on 20000 selected samples for 2000 steps with a learning rate of 1e-6 and a temperature of 1. For each input, 8 responses are generated, with a maximum length of 512 tokens. And the model from the final training step is selected as the final model.
@@ 3.3 Evaluation Metrics
In this experiment, the evaluation metric is WER, which is calculated by extracting the text between <answer>and </answer>tags from the decoded output of the LibriSpeech, Libri2Mix, and Libri3Mix test sets. For the small number of outputs with formatting errors or incomplete tags, the result is treated as empty for evaluation, resulting in deletion errors. The reported WER is computed only for the target speaker. In multi-speaker mixtures, each speaker is independently treated as the target, and the WER is evaluated separately for each speaker.
@@ 4.1 Comparison to Baselines
We compare our proposed model with both traditional TS-ASR methods and LLM-based approaches. As shown in Table 1, our model achieves significant improvements, with WERs of 4.84% on Libri2Mix and 12.23% on Libri3Mix.
Table 1: TS-ASR performance on Libri2Mix and Libri3Mix test-clean dataset. Evaluated by WER(%).
Model LibriMix
2-mix 3-mix
Whisper + LoRA (Ma et al., 2024) 11.98 -
MUSE-TSASR  (Seong et al., 2025) 10.53 -
WhisperTSE-L  (Ma et al., 2025a) 8.1 -
Whisper-SS-TTI (Meng et al., 2024) 7.97 21.97
WavLM + TSE (Huang et al., 2023) 7.6 -
TS-VAD(460h)  (Maeda et al., 2025) 6.61 14.81
MT-LLM  (Meng et al., 2025) 6.7 16.2
TS-ASR Base 7.4 17.24
TS-ASR + CoT 5.29 13.57
TS-ASR + CoT + RL 4.84 12.23
@@
Compared with MT-LLM, the current best-performing LLM-based TS-ASR model, our TS-ASR BASE model without CoT data has a lower performance. However, training with CoT data achieves a significant improvement, reducing the WER on Libri2Mix and Libri3Mix by 21% and 16.2%, respectively. At this stage, the proposed model already outperforms previous work on the LibriMix dataset. It demonstrates that logic reasoning helps the model more accurately transcribe the target speaker.
Based on the model trained on CoT data, RL fine-tuning is applied to selected data, further reducing the WER of 2-mix and 3-mix to 4.84% and 12.23%, respectively. It indicates that RL training can help the model further improve the reasoning capabilities. Specifically, we observe that the rate of incomplete <think>and <answer>tags decreases from 0.4‰  to 0.23‰  after RL fine-tuning, indicating that the format reward further ensures the completeness of the reasoning structure. In terms of final performance, our model achieves substantially lower WER than all baseline models, achieving new SOTA performance in Libri2Mix and Libri3Mix dataset. This confirms that our exploration of LLM-based CoT and RL for TS-ASR represents a valuable and effective direction for research.
Table 2: Comparison of similarity scores and similarity levels for CoT data, evaluated by WER(%).
Model LibriMix
2-mix 3-mix
TS-ASR + CoT
     + similarity score 5.39 14.09
     + similarity level 5.29 13.57
4.2 Effectiveness of Discrete Similarity Levels
We explore the impact of continuous similarity scores and discrete similarity levels on model performance to identify an effective configuration. As shown in Table 2, the model trained on the data using similarity scores is compared to that trained on data with discrete similarity levels. Results show that using discrete similarity levels reduces WER by 0.1% and 0.52%, with consistent improvements observed on both Libri2Mix and Libri3Mix.
The improvement shows the effectiveness of replacing continuous floating-point scores with discrete levels in the CoT data. This further indicates that detailed floating values are more difficult to understand and leverage by LLMs than concise discrete similarity levels.
4.3 Ablation Studies for CoT Template
This section presents ablation studies on the key variables in the CoT template to evaluate the importance of different types of information for TS-ASR reasoning. As shown in Figure 4, the CoT data are processed by individually removing information related to the number of speakers, gender, speaker similarity, and start and end times, followed by retraining the CoT model based on the TS-ASR BASE model.
Figure 4: A sample of CoT data for 2-speaker mixed speech. Different types of removed information are being marked with distinct colors and numbers in the text.
Table 3: Ablation Studies for CoT template, evaluated by WER(%). ”-” means the removal of specific information from the full CoT data.
Model LibriMix
2-mix 3-mix
TS-ASR + full CoT data 5.29 13.57
     - speaker number 5.46 14.01
     - speaker gender 5.32 13.7
     - speaker similarity 6.08 14.93
     - start and end times 5.19 13.54
In Table 3, the results show that removing speaker similarity leads to a significant performance drop. The WER increases from 5.29% to 6.08% on Libri2Mix and from 13.57% to 14.93% on Libri3Mix. This indicates that speaker similarity is the most critical feature for CoT reasoning. Removing the number of speakers also causes a noticeable increase in WER, particularly in the 3-speaker mixture, where WER rises by 0.43%, suggesting that this information is also important. Meanwhile, removing either gender or time information affects performance by no more than 0.2% across all test sets, with no significant difference observed.
This conclusion aligns with human auditory perception. In noisy environments, listeners rely on speaker similarity to identify the target speaker. At the same time, the number of speakers, which reflects the complexity of the acoustic environment, affects how clearly the target speech can be heard. Together, these two factors are essential for speech understanding in cocktail party scenarios.
Figure 5: Schematic illustration of different RL data selection strategies. All strategies select 20,000 samples as RL training data.
4.4 Ablation Studies for RL Data Selection
Since data selection is important for effective RL training, experiments are conducted on different data selection methods for the RL training of TS-ASR task. Table 4 presents the impact of different RL training data selection strategies on model performance. As shown in Figure 5, “Random Sampling” denotes randomly selecting a subset from the full training dataset for RL training. “Balanced Correct&Error” selects correct and error samples in a 1:5 ratio from the model trained on CoT data for RL training, aiming to focus on difficult cases while preserving general performance. The ”Error-only Sampling” strategy uses only the subset of incorrectly predicted samples for subsequent training. “Error-Rate Stratified” sorts all error samples in ascending order of WER and then randomly selects a total of 20,000 samples from the lowest 1/3, middle 1/3, and highest 1/3, with the number of samples drawn from each group following a ratio of 1:6:3.
Table 4: Ablation Studies for RL data selection, evaluated by WER(%).
Model LibriMix
2-mix 3-mix
TS-ASR + CoT (with similarity level) + RL
     + Random Sampling 4.86 12.72
     + Balanced Correct&Error 4.85 12.43
     + Error-Rate Stratified 4.8 12.34
     + Error-only Sampling 4.84 12.23
TS-ASR + CoT (with similarity level - time) + RL
     + Error-Rate Stratified 4.81 12.29
     + Error-only Sampling 4.84 12.32
Experimental results show that the “Random Sampling” strategy performs the worst on Libri3Mix, with a WER of 12.72%, while the “Error-only Sampling” strategy achieves the best result, reducing WER to 12.23%. When considering Libri2Mix and Libri3Mix together, the ”Error-Rate Stratified” strategy also performs well, achieving WERs of 4.80% and 12.34%, respectively. This approach also uses error only samples and further refines sample selection by stratifying them according to WER, thereby focusing more on the challenging cases within the error set. This demonstrates that in reinforcement learning for TS-ASR, challenging error samples contribute more to performance improvement than correct samples. Therefore, training with a higher proportion of error samples is more effective for helping the model correct mistakes and improve overall accuracy.
Since the CoT model without time information achieves marginal better performance in Table 3, we adopt this model and apply the two best data selection strategies for RL training. As shown in the last two rows of Table 4, removing time information does not lead to better final performance compared to using the full CoT data. Therefore, the time information in the CoT data is retained.
Table 5: Ablation Studies for single speaker in LibriSpeech test-clean dataset, evaluated by WER(%).
@@ 4.4 Ablation Studies for RL Data Selection
   TS-ASR Base    8.1
   TS-ASR + CoT    3.88
   TS-ASR + CoT + RL    3.65
4.5 Comparison for Single-speaker
MT-LLM (Meng et al., 2025) reports single-speaker ASR performance alongside its evaluation of multi-talker scenarios under the MT-ASR task. It is also important to report single-speaker performance for the TS-ASR task. In real-world applications, the system cannot know in advance whether the input is a multi-talker mixture or not. As a result, given a reference speech from the target speaker, a TS-ASR system should be required to transcribe an utterance containing only the speech of the target speaker. Consequently, it needs to handle both multi-talker mixtures and single-speaker scenarios. However, prior work has not paid enough attention to single-speaker TS-ASR performance on the LibriSpeech dataset.
To evaluate performance in this setting, ablation studies are conducted on the LibriSpeech test-clean dataset, with results shown in Table 5. Our evaluation follows the standard TS-ASR protocol by using the TS-ASR prompt and randomly selecting a reference utterance from the same speaker. The results show that our proposed improvements enhance TS-ASR performance not only in multi-speaker settings but also in the single-speaker case, with the model trained using CoT and RL achieving a WER of 3.65% on the single-speaker test set.
Nevertheless, TS-ASR still faces unique challenges in single-speaker scenarios. It is observed that in single-speaker scenarios, the model tends to produce more deletion errors. This may be due to the fact that TS-ASR is designed to suppress non-target speakers in multi-talker environments, which can cause the suppression mechanism to mistakenly treat the target speech as interference in single-speaker inputs, leading to degraded performance.
@@ 5 Conclusion
In this work, we proposed a novel framework that incorporates CoT and RL training into the TS-ASR task to enhance reasoning capabilities in cocktail party scenarios. The approach began with training a TS-ASR BASE model using the LALM architecture. Subsequently, a novel methodology was proposed to generate CoT training samples by extracting salient features from overlapped speech, such as number of speakers, speech duration, gender information, and similarity to the reference speech. These features were systematically organized into structured reasoning data to fine-tune the TS-ASR BASE model. Furthermore, a subset of samples from the CoT training phase was selected for additional RL-based refinement. Experimental results showed a significant improvement of TS-ASR performance with CoT and RL training, which not only demonstrated the effectiveness of our constructed CoT dataset and the proposed training framework but also established a SOTA performance in the Libri2Mix and Libri3Mix datasets on the TS-ASR task. The successful integration of CoT and RL into TS-ASR demonstrated the efficacy of logical reasoning in cocktail party scenarios. As LALMs continue to advance with CoT and RL techniques, we believe this framework has significant potential for further solving the cocktail party problem.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.